OpenStudy (anonymous):

Identify the correct slope and y-intercept: of the equation 2x + 3y = 6?

OpenStudy (unklerhaukus):

if you can re arrange to the form y=mx+b the slope will be m and the y-intercept will be b

OpenStudy (anonymous):

and wats the procidure to solve it?

OpenStudy (unklerhaukus):

2x + 3y = 6 minus 2x form both sides, then divide both sides by three

OpenStudy (unklerhaukus):

... y=

OpenStudy (anonymous):

do i divide the 2x by 3 or 6 divided by 3?

OpenStudy (unklerhaukus):

take away 2x first 2x-2x + 3y = 6-2x then divide by three

OpenStudy (anonymous):

i knw i did that already but when i divide do i divide both numbers or jst 2x?

OpenStudy (unklerhaukus):

(6-2x)/3 = 6/3 - 2x/3

OpenStudy (anonymous):

ohh ok that makes sence so the righ answer would be -2/3 y-intercept; (0,2) am i right?

OpenStudy (unklerhaukus):

yes thats right the slope m=-2/3 and b is 2, so the y-intercept is at the point (0,2) ,

OpenStudy (anonymous):

ok thanks